5/ If you put $500 in the bank every month, and earn 2.5% interest, and inflation is an average of 2%, how much will you have in the bank in 10 years, and how much is it worth in today’s money?
Answer #1

First we have to calculate the future value of monthly annuity of $ 500 till 10 years at interest rate of 2.5% per annum.

This is calculated through excel function FV, where,

Rate = 2.5%/12

nper = 10*12 = 120

pmt = 500

type = 0 (i.e at the end of the month)

The answer comes to $ 68085.97. This is the amount that will be received if $ 500 is invested every month for 10 years at interest rate of 2.5%

To calculate the worth of this $ 68085.97 in todays terms, its present value needs to be calculated

PV @ 0.167% for 120 months comes to 0.81887 [ 1 / (1.00167^120)]

Thus PV of 68085.97 = 68085.97 * 0.81887 = $55753.56